# A Survey of Foundation Models for IoT: Taxonomy and Criteria-Based Analysis

### v1

@@ 1. Introduction
The structure of this paper is as follows: Section 2 introduces three fundamental paradigms and frameworks for applying foundation models to IoT tasks, providing essential background for readers new to the field. Sections 3 through 6 examine the four key performance criteria along with commonly used approaches to improve each. Section 7 reviews the evaluation metrics used in the current literature for each criterion and discusses the strategies employed to assess FM-based solutions in IoT applications. Finally, Section 8 discusses existing research limitations and outlines future directions. Figure 1 presents a taxonomy of this paper and representative techniques.
@@
2) [Prompt-Based (§2.1),text width=10.9em [ Xu et al. (2024a); An et al. (2024); Ho et al. (2024); Ouyang and Srivastava (2024); Da et al. (2024); Ji et al. (2024);
Cleland et al. (2024); Song et al. (2023b); Xiao et al. (2024) Chen et al. (2024b); Arrotta et al. (2024); Yang et al. (2024c);
@@
Wang et al. (2024c); Cui et al. (2024b); Huang et al. (2024b); Abdallah et al. (2024); Cui et al. (2024a); Liu et al. (2024d);
Yang et al. (2025) Wen et al. (2024); Xu et al. (2024b); Yang et al. (2024b); Li et al. (2024f); Wang et al. (2024a). ,leaf ,text width=55.5em] ] [Training-Based (§2.3),text width=10.9em [ Ouyang et al. (2025); Mo et al. (2024); Chen et al. (2024a); Xu et al. (2021); Kimura et al. (2024b); Alikhani et al. (2024);
@@
Ouyang and Srivastava (2024); Yang et al. (2025); Sun and Ortiz (2024); Xiao et al. (2024);
@@
Xue et al. (2024); Wan et al. (2024). , leaf, text width= 42.5em] ] [Experience Archive,text width=13.4em [ Cui et al. (2024b); Cui et al. (2024a); Abdallah et al. (2024). ,leaf, text width = 42.5em] ] [Efficient Model Design,text width=13.4em [ Xu et al. (2021); Zhuang et al. (2024). Zhang et al. (2024d). Kimura et al. (2024b);
@@
Awareness (§4) [Data Enrichment,text width=13.4em [ An et al. (2024); Ji et al. (2024). ,leaf, text width=42.5em] ] [Fine-Tuning,text width=13.4em [Guo et al. (2024b). ,leaf, text width= 42.5em] ] [Contextual Inputs,text width=13.4em [Nepal et al. (2024); Cui et al. (2024a); Yang et al. (2025). ,leaf, text width= 42.5em] ] [Contextual Memory,text width=13.4em [Yang et al. (2025). ,leaf, text width= 42.5em] ] [Active Sensing,text width=13.4em [Cui et al. (2024a). ,leaf, text width= 42.5em] ] [Iterative Self-Refinement,text width=13.4em [ Liu et al. (2024b); Cui et al. (2024a); Guo et al. (2024b). ,leaf, text width= 42.5em] ] ] [Safety (§5) [Safety Requirement
Incorporation,text width=13.4em [Xu et al. (2024b). ,leaf, text width= 42.5em] ] [User Request Verification,text width=13.4em [Cui et al. (2024b); Xu et al. (2024b). ,leaf, text width= 42.5em] ] [Post-Response Verification,text width=13.4em [Xu et al. (2024b); Yang et al. (2024b). ,leaf, text width= 42.5em] ] [Interpretable
Decision-Making,text width=13.4em [Cui et al. (2024b). ,leaf, text width= 42.5em] ] [Iterative Self-Refinement,text width=13.4em [Xu et al. (2024b); Yang et al. (2024b). ,leaf, text width= 42.5em] ] ] [Security and
@@
  (§7) [Strategy, text width=4.5em [Groundtruth,text width=9.9em [Kimura et al. (2024b); Yang et al. (2025); Yang et al. (2023a); Xu et al. (2021); Alikhani et al. (2024);
@@
Evaluation,text width=9.9em [Yang et al. (2025); Ouyang and Srivastava (2024); Abdallah et al. (2024); Song et al. (2023b); Kara et al. (2024);
@@
Context-Awareness: Executability rate (Liu et al., 2024b); Personalization score (Yang et al., 2025).
@@
Figure 1. Taxonomy of Foundation Models for IoT. The structure of this paper is as follows: Section 2 introduces three fundamental paradigms and frameworks for applying foundation models to IoT tasks, providing essential background for readers new to the field. Sections 3 through 6 examine the four key performance criteria along with commonly used approaches to improve each. Section 7 reviews the evaluation metrics used in the current literature for each criterion and discusses the strategies employed to assess FM-based solutions in IoT applications. Finally, Section 8 discusses existing research limitations and outlines future directions.
@@ 2.1. Paradigm I: Prompt-Based Methods
Prompt-based methods leverage commercial or open-source pretrained large language models (LLMs) (e.g., GPT-4 (Achiam et al., 2023), LLaMA-2 (Touvron et al., 2023), DeepSeek-V2 (Liu et al., 2024a)) to perform IoT tasks through textual inputs without requiring additional training (Sahoo et al., 2024; Liu et al., 2023). These prompts typically serve as task-specific instructions (e.g., “You are a knowledgeable IoT expert. Please classify the human activity described by the following IMU data.”), which the model interprets to generate a response in the specified format (Sahoo et al., 2024; Liu et al., 2023) (please refer to the example prompt for using LLMs to address the IoT task in Appendix A). Because prompt-based approaches rely solely on pretrained models, they enable rapid prototyping and deployment, particularly in settings with limited computational resources.
@@
Prompt-based methods offer several advantages for applying foundation models to IoT tasks: (1) Efficiency: Unlike training-based methods that require substantial labeled data and computational power to update model parameters (see Section 2.3), prompt-based approaches leverage already trained LLMs which exhibit strong few-shot learning capabilities (Song et al., 2023a; Brown et al., 2020), eliminating the need for fine-tuning or retraining. This significantly reduces deployment costs and time. (2) General Knowledge: LLMs trained on large-scale web data possess extensive general and domain-relevant knowledge, which can be leveraged to interpret IoT data and support tasks (e.g., human activity recognition from IMU signals) (Xu et al., 2024a; An et al., 2024; Ho et al., 2024; Ouyang and Srivastava, 2024; Da et al., 2024; Ji et al., 2024; Cleland et al., 2024; Arrotta et al., 2024; Chen et al., 2024b; Song et al., 2023b). (3) Enhanced Reasoning: Prompting techniques like chain-of-thought (CoT) (Wei et al., 2022) and program-of-thought (PoT) (Chen et al., 2022) can activate commonsense reasoning abilities in LLMs and improve their ability to interpret complex or ambiguous IoT data. This has been shown to significantly boost performance in tasks such as human activity recognition, industrial monitoring, and sensor data analysis (Kök et al., 2024; Xiao et al., 2024). (4) Flexibility with Heterogeneous Data: Prompt-based methods can accommodate both structured (e.g., tabular sensor readings) and unstructured data (e.g., geospatial graphs) by explicitly specifying input and output formats in the prompt (Zong et al., 2025; Fan et al., 2024). This flexibility makes them well-suited to the diverse data types found in IoT systems. (5) Natural Language Interfaces: LLMs can serve as intuitive interfaces for querying or controlling IoT systems using natural language. This lowers the barrier to entry for non-experts and enhances system usability and accessibility. (6) Interpretability: Language-based outputs, especially when coupled with reasoning prompts (e.g., “Explain your answer in a few sentences.”), enhance transparency by revealing the model’s reasoning process. This allows users to verify the outputs and increases trust in the system, unlike many traditional ML models that operate as black boxes (Cui et al., 2024b). (7) Real-Time Insights: Prompted LLMs are capable of processing large streams of sensor data and producing actionable responses in real time, making them suitable for rapid prototyping and deployment in dynamic IoT environments (Zong et al., 2025).
@@ 2.2. Paradigm II: Agent-Based Methods
To overcome the limitations of prompt-based methods, LLM agents (Wang et al., 2024b; Luo et al., 2025; Weng, 2023), which augment pretrained language models with additional modules: external memory, tool integration, and planning capabilities, are gaining increasing adoption across various IoT domains (e.g., smart homes (Rivkin et al., 2024; Yonekura et al., 2024), industrial monitoring (Liang and Tong, 2025; Liu et al., 2024e; Wang and Qin, 2024), autonomous driving (Wang et al., 2024c; Cui et al., 2024b; Huang et al., 2024b)). These agents retain the strengths of prompt-based approaches while addressing their key shortcomings:
@@ 3. Criterion I: Efficiency
To improve the efficiency of foundation models for IoT tasks, four common strategies are identified in the surveyed literature: efficient prompting, experience archive, and efficient model architecture design.
@@
To address the challenge of scaling LLMs to large-scale sensor data, researchers have proposed several strategies to reduce the volume of sensor data included in the prompt while preserving inference efficiency and performance. In the following sections, we review five key approaches that address this issue. For strategies focused on optimizing other components of the prompt (e.g., such as instructions or chain-of-thought (CoT) demonstrations (Wei et al., 2022)), please refer to surveys such as Chang et al. (2024).
@@
Data Summarization. To reduce input volume, raw sensor data can be summarized before being passed to the model. Three common approaches are used for summarizing numerical sensor data. The first relies on external tools (e.g., calculators or Python scripts) to extract high-level statistical features (e.g., mean, variance, or FFT-based metrics) (An et al., 2024). This method is particularly effective when these features are directly relevant to downstream tasks (e.g., an elevated average body temperature may indicate fever, suggesting a potential health concern). The second approach leverages LLMs or lightweight language models (e.g., DistillBert (Sanh et al., 2019)) to generate natural language summaries of raw sensor data (Ouyang and Srivastava, 2024). These models have shown strong capabilities in identifying and articulating key numerical patterns (e.g., trends and state changes) and these summaries are typically much shorter than the original data while preserving essential information. The third approach uses task specific modules (e.g., rule-based methods or lightweight ML models) to extract task related information (e.g., identifying individuals from vibration data or recognizing activities from audio signals) (Yang et al., 2025; Ouyang and Srivastava, 2024; Sun and Ortiz, 2024). Only the extracted outputs are passed to the LLM when these simpler models are insufficient for completing the task (Ouyang and Srivastava, 2024). This approach reduces the LLM’s computational burden, allowing it to focus on high-level reasoning and user interaction.
@@
Experience Archive. The experience archive serves as an external memory (e.g., a database or cache) that stores historical interactions between the LLM-agent and the user, including past plans, tool and executor outputs, error messages from failed executions, and prior contextual information and decisions (Cui et al., 2024b, a; Abdallah et al., 2024). By enabling the agent to retrieve relevant past decisions and responses for similar scenarios or user queries, the experience archive reduces the need for repeated reasoning, planning, or replanning. This significantly improves decision-making efficiency by minimizing redundant computation and inference time.

### v2

@@ 1. Introduction
The structure of this paper is as follows: Section 2 reviews and compares our survey with existing related work. Section 3 introduces three fundamental paradigms and frameworks for applying foundation models to IoT tasks, providing essential background for readers new to the field. Sections 4 through 7 examine the four key performance criteria along with commonly used approaches to improve each. Section 8 reviews the evaluation metrics used in the current literature for each criterion and discusses the strategies employed to assess FM-based solutions in IoT applications. Finally, Section 9 discusses existing research limitations and outlines future directions. Figure 1 presents a taxonomy of this paper and representative techniques.
@@
3) [Prompt-Based (§3.1),text width=10.9em [ Xu et al. (2024a); An et al. (2024); Ho et al. (2024); Ouyang and Srivastava (2024); Da et al. (2024); Ji et al. (2024);
Cleland et al. (2024); Song et al. (2023c); Xiao et al. (2024) Chen et al. (2024b); Arrotta et al. (2024); Yang et al. (2024c);
@@
Wang et al. (2024d); Cui et al. (2024b); Huang et al. (2024b); Abdallah et al. (2024); Cui et al. (2024a); Liu et al. (2024d);
Yang et al. (2025a) Wen et al. (2024); Xu et al. (2024b); Yang et al. (2024b); Li et al. (2024f); Wang et al. (2024a);
Zhong et al. (2024b); Mao et al. (2023); Yang et al. (2025b); Shi et al. (2024b). ,leaf ,text width=55.5em] ] [Training-Based (§3.3),text width=10.9em [ Ouyang et al. (2025); Mo et al. (2024); Chen et al. (2024a); Xu et al. (2021); Kimura et al. (2024b); Alikhani et al. (2024);
@@
Ouyang and Srivastava (2024); Yang et al. (2025a); Sun and Ortiz (2024); Xiao et al. (2024);
@@
Xue et al. (2024); Wan et al. (2024). , leaf, text width= 42.5em] ] [Experience Archive,text width=13.4em [ Cui et al. (2024b); Cui et al. (2024a); Abdallah et al. (2024); Li et al. (2024f); Zhong et al. (2024b);
Shi et al. (2024b); Mao et al. (2023); Shi et al. (2024b). ,leaf, text width = 42.5em] ] [Efficient Model Design,text width=13.4em [ Xu et al. (2021); Zhuang et al. (2024). Zhang et al. (2024d). Kimura et al. (2024b);
@@
Awareness (§5) [Data Enrichment,text width=13.4em [ An et al. (2024); Ji et al. (2024); Li et al. (2024b); Ouyang and Srivastava (2024); Xu et al. (2024a). ,leaf, text width=42.5em] ] [Fine-Tuning,text width=13.4em [Guo et al. (2024b); Tian et al. (2025). ,leaf, text width= 42.5em] ] [Contextual Inputs,text width=13.4em [Nepal et al. (2024); Cui et al. (2024a); Yang et al. (2025a); Tian et al. (2025); Huang et al. (2022);
Song et al. (2023b); Singh et al. (2022). ,leaf, text width= 42.5em] ] [Contextual Memory,text width=13.4em [Yang et al. (2025a); Cui et al. (2024a); Rivkin et al. (2024); Saleh et al. (2025). ,leaf, text width= 42.5em] ] [Active Sensing,text width=13.4em [Cui et al. (2024a); Yang et al. (2025b). ,leaf, text width= 42.5em] ] [Iterative Self-Refinement,text width=13.4em [ Liu et al. (2024b); Cui et al. (2024a); Guo et al. (2024b); Xu et al. (2024b); Liu et al. (2024e);
Yang et al. (2024b). ,leaf, text width= 42.5em] ] ] [Safety (§6) [Safety Requirement
Incorporation,text width=13.4em [Xu et al. (2024b); Ravichandran et al. (2025); Brunke et al. (2025). ,leaf, text width= 42.5em] ] [User Request Verification,text width=13.4em [Cui et al. (2024b); Xu et al. (2024b). ,leaf, text width= 42.5em] ] [Post-Response Verification,text width=13.4em [Xu et al. (2024b); Yang et al. (2024b); Yang et al. (2024d); Quartey et al. (2024);
Ravichandran et al. (2025). ,leaf, text width= 42.5em] ] [Interpretable
Decision-Making,text width=13.4em [Cui et al. (2024b); Zhang et al. (2024b); Nouri et al. (2024). ,leaf, text width= 42.5em] ] [Iterative Self-Refinement,text width=13.4em [Xu et al. (2024b); Yang et al. (2024b); Yang et al. (2024d). ,leaf, text width= 42.5em] ] ] [Security and
@@
  (§8) [Strategy, text width=4.5em [Groundtruth,text width=9.9em [Kimura et al. (2024b); Yang et al. (2025a); Yang et al. (2023a); Xu et al. (2021); Alikhani et al. (2024);
@@
Evaluation,text width=9.9em [Yang et al. (2025a); Ouyang and Srivastava (2024); Abdallah et al. (2024); Song et al. (2023c); Kara et al. (2024);
@@
Context-Awareness: Executability rate (Liu et al., 2024b); Personalization score (Yang et al., 2025a).
@@
Figure 1. Taxonomy of Foundation Models for IoT. The structure of this paper is as follows: Section 2 reviews and compares our survey with existing related work. Section 3 introduces three fundamental paradigms and frameworks for applying foundation models to IoT tasks, providing essential background for readers new to the field. Sections 4 through 7 examine the four key performance criteria along with commonly used approaches to improve each. Section 8 reviews the evaluation metrics used in the current literature for each criterion and discusses the strategies employed to assess FM-based solutions in IoT applications. Finally, Section 9 discusses existing research limitations and outlines future directions.
2. Related Work
Table 1. Comparison with Prior Survey Papers. Unlike previous survey papers that primarily focus on specific IoT application domains (e.g., healthcare, smart homes), this survey introduces a taxonomy centered on performance criteria. We also provide a summary of key foundation model paradigms, a multidimensional comparison of these paradigms, and a detailed overview of evaluation metrics for each performance criterion as well as commonly used evaluation strategies. The content in parentheses reflects the alternative perspectives adopted in prior surveys.
Survey
Papers
Performance Criteria
Perspective
FM Paradigms
& Comparison
Evaluation
Metrics Strategies
Baris et al. (2025) ✘ (Sensor modalities, task types)
Xu et al. (2024c) ✘ (IoT application domains)
Kök et al. (2024) ✘ (IoT application domains, IoT system architecture)
Bhat et al. (2025) ✘ (IoT application domains, IoT system architecture)
Khatiwada et al. (2025) ✘ (IoT application domains)
Sarhaddi et al. (2025) ✘ (IoT application domains)
Our Paper
Existing survey papers on foundation models for IoT have organized their reviews around the following three perspectives (Please also refer to Table 1 for a comparison between prior survey studies and our work.): 1. IoT Application Domains: Nearly all prior surveys (Baris et al., 2025; Xu et al., 2024c; Kök et al., 2024; Bhat et al., 2025; Khatiwada et al., 2025; Sarhaddi et al., 2025) construct their taxonomies based on specific IoT application domains, including robotics, autonomous vehicles, industrial systems, experimental platforms, smart cities, smart homes, smart devices, healthcare, agriculture, communication, education, environmental monitoring, cybersecurity, and privacy. This broad coverage reflects the widespread integration of foundation models across diverse IoT sectors that significantly impact human life. As noted in the introduction, this adoption is largely driven by foundation models’ reduced reliance on labeled data and their ability to generalize across different contexts, which are the advantages over traditional machine learning approaches. 2. IoT System Architecture: Several surveys (Kök et al., 2024; Bhat et al., 2025) examine where foundation models can be deployed within the IoT infrastructure: specifically, at the edge, fog, or cloud layers. These taxonomies align with earlier studies (Zhang and Tao, 2020; Seng et al., 2022; Dou et al., 2023) on the deployment of machine learning models, recognizing foundation models as a subclass of machine learning systems with unique deployment considerations. 3. Sensor Modalities and Task Types: Baris et al. (2025) analyze foundation models in cyber-physical systems (CPS) and IoT along two axes: sensor modalities (e.g., single, multi-modal, and flexible) and task types (e.g., fixed, configurable, selectable, and runtime-specifiable). This focus reflects the nature of IoT tasks, which often involve heterogeneous sensors and require models capable of handling multiple, dynamically defined tasks to ensure generalization and adaptability.
Additionally, two papers specifically address the evaluation of foundation models for IoT. Xu et al. (2024c) summarize commonly used evaluation metrics such as accuracy, RMSE, and success rate. Sarhaddi et al. (2025) introduce the emerging “LLM-as-a-Judge” strategy, where large language models are used to assess the performance of other foundation models or LLMs in IoT tasks, which is an increasingly popular approach following the widespread adoption of LLMs.
Compared to prior surveys, our paper introduces a novel taxonomy based on performance criteria. As discussed in the introduction, identifying shared performance criteria across diverse IoT scenarios (e.g., application domains, deployment locations, sensor types, and task types) enables systematic cross-domain comparison of methods targeting specific performance goals. This approach also supports the informed selection of appropriate techniques for new scenarios and fosters the development of more generalizable methods. These insights and the associated research opportunities have been largely overlooked in previous surveys.
Furthermore, to facilitate comparison and selection of methods for each performance criterion, our survey summarizes a broader set of evaluation metrics for each criterion beyond general metrics outlined in Xu et al. (2024c). We also provide a comprehensive discussion of four evaluation strategies: ground-truth comparison, human evaluation, and LLM-as-a-Judge, including their advantages, limitations, and applicable performance criteria. This analysis offers significantly greater depth and coverage than that presented in Sarhaddi et al. (2025).
We also summarize three fundamental paradigms of foundation models for IoT tasks (i.e., prompt-based, agent-based, and training-based methods), which have been overlooked in previous surveys. Understanding these paradigms is crucial because they offer distinct trade-offs in terms of computational efficiency, adaptability to diverse IoT environments, ease of integration with external systems, and requirements for training data and infrastructure. In addition, by clarifying the strengths and limitations of each approach as well as presenting a decision tree to help select the most suitable paradigm, our survey equips researchers and practitioners with the necessary framework to select, adapt, or develop solutions that are best suited to their specific IoT applications and resource constraints.
@@ 3.1. Paradigm I: Prompt-Based Methods
Prompt-based methods leverage commercial or open-source pretrained large language models (LLMs) (e.g., GPT-4 (Achiam et al., 2023), LLaMA-2 (Touvron et al., 2023), DeepSeek-V2 (Liu et al., 2024a)) to perform IoT tasks through textual inputs without requiring additional training (Sahoo et al., 2024; Liu et al., 2023b). These prompts typically serve as task-specific instructions (e.g., “You are a knowledgeable IoT expert. Please classify the human activity described by the following IMU data.”), which the model interprets to generate a response in the specified format (Sahoo et al., 2024; Liu et al., 2023b) (please refer to the example prompt for using LLMs to address the IoT task in Appendix A). Because prompt-based approaches rely solely on pretrained models, they enable rapid prototyping and deployment, particularly in settings with limited computational resources.
@@
Prompt-based methods offer several advantages for applying foundation models to IoT tasks: (1) Efficiency: Unlike training-based methods that require substantial labeled data and computational power to update model parameters (see Section 3.3), prompt-based approaches leverage already trained LLMs which exhibit strong few-shot learning capabilities (Song et al., 2023a; Brown et al., 2020), eliminating the need for fine-tuning or retraining. This significantly reduces deployment costs and time. (2) General Knowledge: LLMs trained on large-scale web data possess extensive general and domain-relevant knowledge, which can be leveraged to interpret IoT data and support tasks (e.g., human activity recognition from IMU signals) (Xu et al., 2024a; An et al., 2024; Ho et al., 2024; Ouyang and Srivastava, 2024; Da et al., 2024; Ji et al., 2024; Cleland et al., 2024; Arrotta et al., 2024; Chen et al., 2024b; Song et al., 2023c). (3) Enhanced Reasoning: Prompting techniques like chain-of-thought (CoT) (Wei et al., 2022) and program-of-thought (PoT) (Chen et al., 2022) can activate commonsense reasoning abilities in LLMs and improve their ability to interpret complex or ambiguous IoT data. This has been shown to significantly boost performance in tasks such as human activity recognition, industrial monitoring, and sensor data analysis (Kök et al., 2024; Xiao et al., 2024). (4) Flexibility with Heterogeneous Data: Prompt-based methods can accommodate both structured (e.g., tabular sensor readings) and unstructured data (e.g., geospatial graphs) by explicitly specifying input and output formats in the prompt (Zong et al., 2025; Fan et al., 2024). This flexibility makes them well-suited to the diverse data types found in IoT systems. (5) Natural Language Interfaces: LLMs can serve as intuitive interfaces for querying or controlling IoT systems using natural language. This lowers the barrier to entry for non-experts and enhances system usability and accessibility. (6) Interpretability: Language-based outputs, especially when coupled with reasoning prompts (e.g., “Explain your answer in a few sentences.”), enhance transparency by revealing the model’s reasoning process. This allows users to verify the outputs and increases trust in the system, unlike many traditional ML models that operate as black boxes (Cui et al., 2024b). (7) Real-Time Insights: Prompted LLMs are capable of processing large streams of sensor data and producing actionable responses in real time, making them suitable for rapid prototyping and deployment in dynamic IoT environments (Zong et al., 2025).
@@ 3.2. Paradigm II: Agent-Based Methods
To overcome the limitations of prompt-based methods, LLM agents (Wang et al., 2024b; Luo et al., 2025; Weng, 2023), which augment pretrained language models with additional modules: external memory, tool integration, and planning capabilities, are gaining increasing adoption across various IoT domains (e.g., smart homes (Rivkin et al., 2024; Yonekura et al., 2024), industrial monitoring (Liang and Tong, 2025; Liu et al., 2024e; Wang and Qin, 2024), autonomous driving (Wang et al., 2024d; Cui et al., 2024b; Huang et al., 2024b)). These agents retain the strengths of prompt-based approaches while addressing their key shortcomings:
@@ 4. Criterion I: Efficiency
To improve the efficiency of foundation models for IoT tasks, four common strategies are identified in the surveyed literature: efficient prompting, experience archive, and efficient model architecture design. Table 3 provides an overview of these methods, including their descriptions and the paradigms (prompt-based, agent-based, training-based) to which they are mostly commonly applied.
Table 3. Summary of Efficiency-Enhancing Methods, with Descriptions and Typically Applicable Paradigms (Prompt-Based, Agent-Based, Training-Based).
Typically Applicable Paradigm
Method Description Prompt Agent Training
Efficient Prompting
Minimizing prompt length
(especially sensor signals) without significantly
compromising the critical information.
Experience Archive
Using long-term external memory to store
task-specific past experiences
(context, user query, LLM agent response).
Efficiency Model Architecture Design
Designing efficient architecture
to reduce computational requirements.
@@
To address the challenge of scaling LLMs to large-scale sensor data, researchers have proposed several strategies to reduce the volume of sensor data included in the prompt while preserving inference efficiency and performance. In the following sections, we review five key approaches that address this issue. In the following sections, we review four key approaches commonly used in prompt-based methods (Section 3.1) and one approach frequently applied in training-based methods (Section 3.3). These efficient prompting techniques are generally not adopted in agent-based methods (Section 3.2), as sensor data are typically preprocessed by external tools or lightweight models before being passed to the LLM (Cui et al., 2024a; Zhong et al., 2024b; Li et al., 2024f; Cui et al., 2024b; Mao et al., 2023; Yang et al., 2025b). For strategies focused on optimizing other components of the prompt (e.g., such as instructions or chain-of-thought (CoT) demonstrations (Wei et al., 2022)), please refer to surveys such as Chang et al. (2024).
@@
Figure 2. Illustration of data summarization methods for enhancing the efficiency of foundation models in IoT tasks, exemplified by an LLM-based Just-in-Time Adaptive Intervention (JITAI) system in mobile health (Nahum-Shani et al., 2016).
Data Summarization. To reduce input volume, raw sensor data can be summarized before being passed to the model. Three common approaches are used for summarizing numerical sensor data. The first relies on external tools (e.g., calculators or Python scripts) to extract high-level statistical features (e.g., mean, variance, or FFT-based metrics) (An et al., 2024). This method is particularly effective when these features are directly relevant to downstream tasks (e.g., an elevated average body temperature may indicate fever, suggesting a potential health concern). The second approach leverages LLMs or lightweight language models (e.g., DistillBert (Sanh et al., 2019)) to generate natural language summaries of raw sensor data (Ouyang and Srivastava, 2024). These models have shown strong capabilities in identifying and articulating key numerical patterns (e.g., trends and state changes) and these summaries are typically much shorter than the original data while preserving essential information. The third approach uses task specific modules (e.g., rule-based methods or lightweight ML models) to extract task related information (e.g., identifying individuals from vibration data or recognizing activities from audio signals) (Yang et al., 2025a; Ouyang and Srivastava, 2024; Sun and Ortiz, 2024). Only the extracted outputs are passed to the LLM when these simpler models are insufficient for completing the task (Ouyang and Srivastava, 2024). This approach reduces the LLM’s computational burden, allowing it to focus on high-level reasoning and user interaction. Figure 2 further illustrates aforementioned methods for enhanced clarity.
@@
Experience Archive. The Experience Archive serves as a form of long-term external memory, in contrast to short-term memory used for in-context learning and internal memory encoded in model parameters (Abdallah et al., 2024; Cui et al., 2024a; Li et al., 2024f; Zhong et al., 2024b). It stores task-specific experiences as a sequence of tuples $\mathcal{H}_{t_{m}:t_{n}}=\{(c_{i},q_{i},r_{i})|i=t_{m}:t_{n}\}$, where $c_{i}$ denotes the context, $q_{i}$ represents the user query or instruction, and $r_{i}$ includes the LLM agent’s response along with the execution result produced by an external tool or executor. For example, $r_{i}$ may include a planned trajectory generated for an autonomous driving task based on prior environmental states and user input (Cui et al., 2024b; Mao et al., 2023). This type of memory is also referred to as episodic memory in prior work (Xiong et al., 2025; Nuxoll and Laird, 2007).
In Section 5, we introduce another form of long-term external memory, referred to as the context repository or semantic memory, which stores knowledge and information about the agent’s current context (e.g., environmental or personal state). The key distinction from the experience archive lies in temporal scope and structure: the experience archive captures past contextual information, interaction history, and execution results, whereas the context repository contains current factual knowledge used for reasoning and decision-making. These facts may be extracted from sensors (e.g., objects detected in a room), pre-stored by developers (e.g., maps or traffic regulations), or inferred from past experiences (e.g., personality traits derived from previous conversations). However, unlike past experiences stored in the experience archive, they do not include the original background data, the LLM agent’s responses, or the corresponding execution results from which they were derived (Nuxoll and Laird, 2007).
The experience archive can significantly improve the efficiency of LLM-based agents in IoT applications. When the current task closely matches a previously encountered one, based on contextual information and user instructions, the agent can retrieve and reuse the prior response as output, avoiding the need for regeneration, which is typically more time-consuming than retrieval (Cui et al., 2024b; Abdallah et al., 2024; Li et al., 2024f; Zhong et al., 2024b). In addition to successful outcomes, storing failed experiences also contributes to efficiency: it enables self-reflection and self-improvement by helping the agent recognize and avoid repeating past mistakes, thereby conserving computational resources (Cui et al., 2024a; Abdallah et al., 2024). Moreover, even when no exact match is found, partially similar past experiences can serve as in-context demonstrations, reducing the need for manual prompt engineering and further streamlining the decision-making process (Xiong et al., 2025).
The experience archive can include both experiences generated during interactions between the LLM agent and the user, as well as preloaded experiences (e.g., from training or validation datasets (Li et al., 2024f; Xiong et al., 2025)). These experiences may be represented in textual form (e.g., natural language (Mao et al., 2023), programming code (Shi et al., 2024b)) or as vector embeddings, similar to those used in retrieval-augmented generation (RAG) systems (Lewis et al., 2020).
To retrieve relevant experiences, similarity metrics are applied (e.g., cosine similarity for embeddings or LLM-defined semantic similarity for text-based comparisons). This process involves interpreting the semantic content of user instructions. However, user inputs may be under-specified or embedded in diverse linguistic patterns, and the LLM may require multiple interaction rounds to fully understand the task. To address this, systems like LLMind (Cui et al., 2024a) prompt the LLM to generate concise summaries of user instructions for more accurate retrieval.
Effective use of the experience archive also depends on proper memory management (Xu et al., 2025; Zhang et al., 2024a; Xiong et al., 2025). First, external memory is inherently limited, making it infeasible to store all past experiences. Second, not all historical experiences are equally valuable; outdated or erroneous memories may mislead the agent. To mitigate these issues, several memory management strategies have been proposed, including selective addition and deletion (Xiong et al., 2025; Zhong et al., 2024b), structural transformation (Wang et al., 2023; Anokhin et al., 2024; Zeng et al., 2024), merging (Zhong et al., 2024a; Liu et al., 2023a; Hu et al., 2024), summarization (Wang et al., 2024c; Zhong et al., 2024a), and reflection. Readers are referred to these works for further details.